if x/(y+z) = y/(z+x) =z/(x+y) then prove (x+y+z)=0 or x=y=z

if  x/(y+z) = y/(z+x) =z/(x+y)  then prove (x+y+z)=0 or x=y=z

Grade:10

1 Answers

Arun
6 years ago
Dear Student
 
let us assume
x+y+z = 0 is true then we have to prove the given part
 
now
y+z = –x
x+y = –z
x+z= –y
 
now 
x/ –x = – 1
y/ –y = – 1
z/ –z = – 1
 
hence proved
 
similarily
by x= y =z you can also prove 
 
Regards
Arun (askIITians forum expert)